OpenStudy (anonymous):

When using the substitution method, what expression for y in the first equation would be substituted in the second equation? 5x – y = –2 9x + 2y = 23

OpenStudy (anonymous):

–5x + 2 / 5x + 2 / –5x – 2 / 5x – 2

OpenStudy (anonymous):

it would be y = 2 +5x because when you solve for y (in the first equation), you have to subtract 5x and when we do that we get... -y = -2 - 5x we can't have a negative variable so we would multiply everything by -1 to get rid of the negative (negative times negative equals a positive) y = 2 + 5x

OpenStudy (anonymous):

y = 5x + 2

OpenStudy (anonymous):

yay you explain it too~ thanks :)

OpenStudy (anonymous):

NP :) watch this video before anything else http://www.khanacademy.org/video/systems-of-equations?playlist=Algebra

OpenStudy (anonymous):

i did! lol. i just don't like math... i am acually doing my sisters work, believe it or not. and i am just getting tired of doing nothing but math prolbem after math problem~

OpenStudy (anonymous):

Hey your sister should do her own homework!